OpenStudy (anonymous):

Utilizing dimensional analysis, find the number of inches in a kilometer given that 1 inch = 2.54 centimeters.

OpenStudy (kirbykirby):

\[ x \text{ inches}=1 \cancel{\text{ km}}\times \frac{100,000 \cancel{\text{ cm}}}{1 \cancel{\text{ km}}}\times \frac{1 \text { inches}}{2.54 \cancel{\text { cm}}}=39,370.079 \text { inches}\]